# HelmetPoser: A Helmet-Mounted IMU Dataset for Data-Driven Estimation of Human Head Motion in Diverse Conditions

### v1

@@ Abstract
footnotetext: This research is supported by China-Singapore International Joint Research Institute (CSIJRI) Development Plan for the Technology Application Center (TAC) and the National Research Foundation, Singapore, under its Medium-Sized Center for Advanced Robotics Technology Innovation (CARTIN). All authors are with School of Electrical and Electronic Engineering, Nanyang Technological University, Singapore 639798, 50 Nanyang Avenue. Kun Cao is with the Department of Control Science and Engineering, College of Electronics and Information Engineering, and Shanghai Research Institute for Intelligent Autonomous Systems, Tongji University, Shanghai, China. (E-mail: jianping.li@ntu.edu.sg, leng0039@e.ntu.edu.sg, jinxin.liu@ntu.edu.sg, xinhang.xu@ntu.edu.sg, tongxing.jin@ntu.edu.sg, mqcao@ntu.edu.sg, thienminh.nguyen@ntu.edu.sg, shyuan@ntu.edu.sg, caokun@tongji.edu.cn, elhxie@ntu.edu.sg) (Jianping Li and Qiutong Leng are co-first authors.)
1 Introduction
@@
In recent years, various data-driven methods have been proposed to address the error accumulation problem in Inertial Measurement Units (IMUs), such as correcting IMU biases [12] or motion drift [13, 14] through deep learning techniques. However, most current research relies on datasets collected in specific platforms, like vehicles [15], robotic platforms [16], and the human body [17]. There are still few datasets that capture the high-frequency motion patterns of the human head, as experienced with helmet usage. The motion patterns of the human head are highly non-linear and uncertain [8], complicating the potential of data-driven IMU motion estimation. This gap presents a significant challenge for achieving high-precision localization systems using the data-driven IMU state estimation on the helmets.
@@
  1. 1.
    We introduce a publicly available dataset that captures human head motion patterns using a helmet-mounted system. This dataset includes motion data from five males and five females, covering a range of daily activities such as walking, running, and stair climbing. It can be used to advance and validate various localization methods.
  2. 2.
    We demonstrated the feasibility of using neural networks to learn IMU biases on helmet-mounted localization devices. Specifically, we employed LSTM and Transformer networks and tested them on the dataset to evaluate their effectiveness.
  3. 3.
    We assessed the performance of the proposed method by varying the dimensions of the IMU data window, different motion patterns, and various IMU sensors. We provided evaluation metrics to establish a baseline for helmet-based localization devices, thus supporting future research in this domain.
@@ 2.1 Sensor Setup of the Helmet Positioning System
2.2 IMU Pre-integration for Ground Truth IMU Bias Estimation
@@ 2.2.2 GT IMU Bias Calculation
$\begin{bmatrix}\alpha_{b_{k}+1}^{b_{k}}\\ \beta_{b_{k}+1}^{b_{k}}\\ \gamma_{b_{k}+1}^{b_{k}}\\ 0\\ 0\end{bmatrix}=\begin{bmatrix}\mathbf{R}_{w}^{b_{k}}(\mathbf{p}_{b_{k}+1}^{w}-\mathbf{p}_{b_{k}}^{w}+\frac{1}{2}\mathbf{g}^{w}\Delta t_{k}^{2}-\mathbf{v}_{b_{k}}^{w}\Delta t_{k})\\ \mathbf{R}_{w}^{b_{k}}(\mathbf{v}_{b_{k}+1}^{w}+\mathbf{g}^{w}\Delta t_{k}-\mathbf{v}_{b_{k}}^{w})\\ \mathbf{q}_{b_{k}}^{w^{-1}}\otimes\mathbf{q}_{b_{k}+1}^{w}\\ \mathbf{b}_{a_{b_{k}+1}}-\mathbf{b}_{a_{b_{k}}}\\ \mathbf{b}_{w_{b_{k}+1}}-\mathbf{b}_{w_{b_{k}}}\end{bmatrix},$ (9)

### v2

@@ Abstract
footnotetext: This research is supported by China-Singapore International Joint Research Institute (CSIJRI) Development Plan for the Technology Application Center (TAC) and the National Research Foundation, Singapore, under its Medium-Sized Center for Advanced Robotics Technology Innovation (CARTIN). All authors are with China-Singapore International Joint Research Institute (CSIJRI) and School of Electrical and Electronic Engineering, Nanyang Technological University, Singapore 639798, 50 Nanyang Avenue. Kun Cao is with the Department of Control Science and Engineering, College of Electronics and Information Engineering, and Shanghai Research Institute for Intelligent Autonomous Systems, Tongji University, Shanghai, China. (E-mail: jianping.li@ntu.edu.sg, leng0039@e.ntu.edu.sg, jinxin.liu@ntu.edu.sg, xinhang.xu@ntu.edu.sg, tongxing.jin@ntu.edu.sg, mqcao@ntu.edu.sg, thienminh.nguyen@ntu.edu.sg, shyuan@ntu.edu.sg, caokun@tongji.edu.cn, elhxie@ntu.edu.sg) (Jianping Li and Qiutong Leng are co-first authors.)
1 Introduction
@@
In recent years, various data-driven methods have been proposed to address the error accumulation problem in Inertial Measurement Units (IMUs), such as correcting IMU biases [12] or motion drift [13, 14] through deep learning techniques. However, most current research relies on data sets collected on specific platforms, such as vehicles [15], robotic platforms [16, 17, 18], and the human body [19]. There are still few datasets that capture the high-frequency motion patterns of the human head, as experienced with helmet usage. The motion patterns of the human head are highly non-linear and uncertain [8], complicating the potential of data-driven IMU motion estimation. This gap presents a significant challenge for achieving high-precision localization systems using the data-driven IMU state estimation on the helmets.
@@
(1) We introduce a publicly available dataset that captures human head motion patterns using a helmet-mounted system. This dataset includes motion data from five males and five females, covering a range of daily activities such as walking, running, and stair climbing. It can be used to advance and validate various localization methods.
(2) We demonstrated the feasibility of using neural networks to learn IMU biases on helmet-mounted localization devices. Specifically, we employed LSTM and Transformer networks and tested them on the dataset to evaluate their effectiveness.
(3) We assessed the performance of the proposed method by varying the dimensions of the IMU data window, different motion patterns, and various IMU sensors. We provided evaluation metrics to establish a baseline for helmet-based localization devices, thus supporting future research in this domain.
@@ 2.1 Sensor Setup of the Helmet Positioning System
2.2 Pre-integration for Ground Truth IMU Bias Estimation
@@ 2.2.2 GT IMU Bias Calculation
$\small\begin{bmatrix}\alpha_{b_{k}+1}^{b_{k}}\\ \beta_{b_{k}+1}^{b_{k}}\\ \gamma_{b_{k}+1}^{b_{k}}\\ 0\\ 0\end{bmatrix}=\begin{bmatrix}\mathbf{R}_{w}^{b_{k}}(\mathbf{p}_{b_{k}+1}^{w}-\mathbf{p}_{b_{k}}^{w}+\frac{1}{2}\mathbf{g}^{w}\Delta t_{k}^{2}-\mathbf{v}_{b_{k}}^{w}\Delta t_{k})\\ \mathbf{R}_{w}^{b_{k}}(\mathbf{v}_{b_{k}+1}^{w}+\mathbf{g}^{w}\Delta t_{k}-\mathbf{v}_{b_{k}}^{w})\\ \mathbf{q}_{b_{k}}^{w^{-1}}\otimes\mathbf{q}_{b_{k}+1}^{w}\\ \mathbf{b}_{a_{b_{k}+1}}-\mathbf{b}_{a_{b_{k}}}\\ \mathbf{b}_{w_{b_{k}+1}}-\mathbf{b}_{w_{b_{k}}}\end{bmatrix},$ (9)